OpenStudy (anonymous):

For f(x) = 5 – 3x and g(x) = x^2 – 1, find f(g(x)) and g(f(x)).

OpenStudy (anonymous):

f(g(x)) Just means that put in g(x) for x in the function f(x). 5-3(x^2-1) 5-3x^2-3 3x^2+2

OpenStudy (anonymous):

thanks for the help! :)

OpenStudy (anonymous):

No problem. Do you understand how to do the next one?

OpenStudy (jdoe0001):

$$ f(x) = \color{green}{5 – 3x}\\ g(x) = \color{blue}{x^2 – 1}\\ f(g(x))=5 – 3(\color{blue}{x^2 – 1})\\ g(f(x))=(\color{green}{5 – 3x})^2 – 1 $$

OpenStudy (anonymous):

Yes!